Solve: 5/4x + 1/8x = 7/8x + x
Answer by stanbon(48558)

Solve: (5/4)x + (1/8)x = (7/8)x + x
---------------------
(11/8)x = (15/8)x
=======================
That is a contradiction.
No solution.
==============================
Cheers,
Stan H>